OpenStudy (anonymous):

Please help me double check functions. f(x)=4x-1 h(x)=-x+3 (FoH)(7) = -17??

OpenStudy (anonymous):

F(h(7)) = 4(-7+3)-1 =4(-4)-1 =-16-1 =-17 That would be correct.

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

You're welcome :)